Using the circle graph that you have made, formulate at least two problems involving arcs and central angles, then solve.


1
Expert's answer
2021-02-24T06:54:49-0500

The different school fees of a typical and ordinary student like me are represented in a circle graph above namely: “School Fees”. Let say that I have Php 5000 budget for different school fees for the whole school year. I spend, Php 2000 for extra-curricular activities in different subject area, Php 1500 for research fund, Php 500 for membership fees of different clubs, Php 200 for photocopy of paper works, and Php 800 for school projects.

a). In the circle graph, what is the measure of the central angle corresponding to each item?

By following the formula: (amount of money per school fees ÷ budget for the

whole school year-Php 5000) x 360

Extra-curricular activities in different subject area

("\frac{2000}{5000}" ) x 360 = 144°

Research fund

("\frac{1500}{5000}" ) x 360 = 108°

Membership fees of different clubs

("\frac{500}{5000}" ) x 360 = 36°

Photocopy of paper works

("\frac{200}{5000}" ) x 360 = 14°

School projects

("\frac{800}{5000}" ) x 360 = 58°

b. Suppose the radius of the circle graph is 35 cm. What is the area of each sector in the

circle graph?

Extra-curricular activities in different subject area

= 1538.6

Research fund

= 1153.95

Membership fees of different clubs

= 384.65

Photocopy of paper works

= 149.58611

School projects

= 619.71388